OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x^4 + 36x^2

hartnn (hartnn):

factor out x^2, what u get ?

OpenStudy (anonymous):

x^2 + 36 ?

hartnn (hartnn):

yes, x^2(x^2+36) now u know how u can factor x^2+36 ?

OpenStudy (anonymous):

(x + 6) (x + 6) ?

hartnn (hartnn):

nopes, if it would be x^2-36 , then its (x + 6) (x + 6) are u aware of complex numbers or i ?

OpenStudy (anonymous):

but postive times positve is postive. I understand though. I am but i'm a bit confused.

OpenStudy (anonymous):

i know theres an i in the answer, i just forgot how to get there

hartnn (hartnn):

ok, x^2 + a^2 = (x+ai)(x-ai)

hartnn (hartnn):

now? write 36 as ?

OpenStudy (anonymous):

so x+6i (x-6i) ?

hartnn (hartnn):

careful with brackets ( x+6i) (x-6i)

OpenStudy (anonymous):

so the answer is f(x) = x^2(x + 6i)(x - 6i)

hartnn (hartnn):

that is correct

OpenStudy (anonymous):

thank you! I have three more for today! Youre awesome!

hartnn (hartnn):

welcome ^_^